OpenStudy (anonymous):

A 10−gallon container of water costs (y + 2) dollars while a 5−gallon container of water costs (y − 2) dollars. If the price per gallon of water is the same for both containers, the equation below can be used to solve for y. y+2/10=y-2/5 What is the value, in dollars, of y? (Do not put a $ in your answer.) (4 points)

OpenStudy (anonymous):

help mi please

OpenStudy (binary3i):

equation that you have written is wrong. (y+2)/10 = (y-2)/5 y = 6